# Periodic orbits of non-degenerate lacunary contact forms on prequantization bundles

### v2

@@ 1.2. Organization of the paper
The rest of this paper is organized as follows. Sections 2 and 3 furnish the necessary background for the proof of Theorem 1.6. More precisely, in Section 2.1 we briefly review several facts about positive equivariant symplectic homology. This homology is computed for suitable prequantization $S^{1}$-bundles in Section 2.2. The resonance relations for equivariant symplectic homology are presented in Section 2.3. Section 3 states the index recurrence theorem, which is a key tool in the proof of Theorem 1.6. The proofs of Theorems 1.6 and 1.13 are established in Sections 4 and 5 respectively. Section 6 presents results about the multiplicity of periodic Reeb orbits on lens spaces and their unit cosphere bundles, extending some results from [27]. Finally, Section 7 poses some questions concerning contact forms with finitely many closed orbits.
@@ 2.1. Equivariant symplectic homology
and an action filtration introduced by McLean and Ritter [41]; cf. [1, Section 2]. Let $\alpha$ be a non-degenerate contact form on $M$ supporting the contact structure $\xi$. Recall that a periodic orbit $\gamma$ of $\alpha$ is good if its index has the same parity of the index of the underlying simple closed orbit. Then the positive equivariant symplectic homology ${\mathrm{SH}}_{*}^{S^{1},+}(W)$ with coefficients in ${\Lambda}$ is the homology of a complex ${\mathrm{CC}}_{*}(\alpha)$ generated by the good closed Reeb orbits of $\alpha$; see [24, Proposition 3.3]. (More precisely, [24, Proposition 3.3] is proved assuming that $\Omega|_{\pi_{2}(W)}=0$ and using $\mathbb{Q}$-coefficients but its proof can be readily adapted to our context since it is purely algebraic; cf. [1, Section 2].) This complex is graded by the Conley–Zehnder index and filtered by the action. Furthermore, once we fix a free homotopy class of loops in $W$, the part of ${\mathrm{CC}}_{*}(\alpha)$ generated by closed Reeb orbits in that class is a subcomplex. As a consequence, the entire complex ${\mathrm{CC}}_{*}(\alpha)$ breaks down into a direct sum of such subcomplexes indexed by free homotopy classes of loops in $W$.
@@
This description of the positive equivariant symplectic homology as the homology of ${\mathrm{CC}}_{*}(\alpha)$ is not quite standard, but it is most suitable for our purposes. (We refer the reader to [24] for more details and further references and to [8, 9, 10, 11, 12, 42] for the original construction of the equivariant symplectic homology.) To see why ${\mathrm{HC}}_{*}(M):={\mathrm{SH}}_{*}^{S^{1},+}(W)$ can be obtained as the homology of a single complex generated by good closed Reeb orbits, let us first consider an admissible Hamiltonian $H$ on the symplectic completion of $W$ and focus on the orbits of $H$ with positive action. Such orbits are in a one-to-one correspondence with closed Reeb orbits $\gamma$ with action below a certain threshold $T$ depending on the slope of $H$. The $S^{1}$-equivariant Floer homology of $H$ is the homology of a Floer-type complex obtained from a non-degenerate parametrized perturbation of $H$; [11, 42]. This complex is filtered by the action. (Here we are using the action filtration introduced by McLean-Ritter [41].) The $E^{1}$-term of the resulting spectral sequence (over ${\Lambda}$) is generated by the good Reeb orbits of $\alpha$ with action below $T$. Now we can (canonically, once the generators are fixed) reassemble the differentials $\partial_{r}$ into a single differential $\partial$ on ${\mathrm{CC}}_{*}(H):=E^{1}_{*,*}$ in such a way the the homology of the resulting complex is $E^{\infty}={\mathrm{HF}}_{*}^{S^{1},+}(H)$. Roughly speaking, $\partial=\partial_{1}+\partial_{2}+\ldots$, where $\partial_{r}$ is suitably “extended” from $E^{r}$ to $E^{1}$. Moreover, this procedure respects the action filtration and is functorial with respect to continuation maps. Passing to the limit in $H$, we obtain the complex ${\mathrm{CC}}_{*}(\alpha)$ as the limit of the complexes ${\mathrm{CC}}_{*}(H)$; see [24, Sections 2.5 and 3] for further details.
[…]
@@ 2.2. Equivariant symplectic homology of prequantizations
The next proposition, essentially taken from [27, Proposition 3.1], shows how to compute the equivariant symplectic homology of a suitable prequantization in terms of the homology of the basis. This computation will be crucial throughout this work.
@@
    ${\mathrm{HC}}^{0}_{\ast}(M)\cong\bigoplus_{m\in\mathbb{N}}{\mathrm{H}}_{\ast-2mc_{B}+n}(B;{\Lambda}).$ (2.1)
@@
    ${\mathrm{HC}}^{0}_{\ast}(M)\cong\bigoplus_{m\in\mathbb{N}}{\mathrm{H}}_{\ast+2mc_{B}-n}(B;{\Lambda}).$ (2.2)
@@
    Alternatively, assume that $B$ is spherically positive monotone with $c_{B}\geq 2$ and, as in the hypothesis (NF) of Theorem 1.6, $c_{1}(\xi)|_{H_{2}(M,\mathbb{Q})}=0$ and $\alpha$ is a non-degenerate contact form on $(M,\xi)$ such that all contractible closed Reeb orbits have index greater than $3-n$. Then (2.1) also holds.
@@
Remark 2.3.
Note that the requirement that $c_{B}\geq 2$ from (b) is automatically satisfied in the setting of Theorem 1.6 as a consequence of the assumptions $\omega|_{\pi_{2}(B)}\neq 0$ and $c_{B}>n/2$. Indeed, $c_{B}\geq 2$ when $n>1$ and for $n=1$ we necessarily have $B=S^{2}$ and hence $c_{B}=2$.
@@
First note that, by the universal coefficient theorem, ${\mathrm{H}}_{*}(B;{\Lambda})\cong{\mathrm{H}}_{*}(B;\mathbb{Q})\otimes{\Lambda}$ since ${\Lambda}$ is a field. Then the proof goes word-for-word the proof of [27, Proposition 3.1] using the Novikov field and the action filtration mentioned in the previous section, except in the proof of [27, Lemma 3.3], used in the proof of item (b), where it is assumed that $c_{1}(\xi)=0$ while here we are allowing $c_{1}(\xi)$ to be torsion. This lemma establishes that a sufficiently small non-degenerate perturbation of the connection form is index-admissible, that is, every contractible periodic orbit $\gamma$ satisfies ${\mu}(\gamma)>3-n$. The assumption in [27, Lemma 3.3] that $c_{1}(\xi)=0$ is used to ensure that the determinant line bundle $\Lambda^{n}_{\mathbb{C}}\xi$ is trivial. When $c_{1}(\xi)$ is torsion it is no longer true in general, but we have that $(\Lambda_{\mathbb{C}}^{n}\xi)^{\otimes N}$ is a trivial line bundle, where $N$ is the smallest positive integer such that $Nc_{1}(\xi)=0$. Choose a trivialization $\tau:(\Lambda_{\mathbb{C}}^{n}\xi)^{\otimes N}\to M\times\mathbb{C}$ which corresponds to a choice of a non-vanishing section ${\mathfrak{s}}$ of $(\Lambda_{\mathbb{C}}^{n}\xi)^{\otimes N}$. The choice of this trivialization furnishes a unique way to symplectically trivialize $\oplus_{1}^{N}\xi$ along periodic orbits of $\alpha$ up to homotopy. As a matter of fact, given a periodic orbit $\gamma$, let $\Phi:{\gamma^{*}\oplus_{1}^{N}\xi}\to S^{1}\times\mathbb{C}^{nN}$ be a trivialization of $\oplus_{1}^{N}\xi$ over $\gamma$ as a Hermitian vector bundle such that its highest complex exterior power coincides with $\tau$. This condition fixes the homotopy class of $\Phi$: given any other such trivialization $\Psi$ we have, for every $t\in S^{1}$, that $\Phi_{t}\circ\Psi_{t}^{-1}:\mathbb{C}^{nN}\to\mathbb{C}^{nN}$ has complex determinant equal to one and therefore the Maslov index of the symplectic path $t\mapsto\Phi_{t}\circ\Psi_{t}^{-1}$ vanishes, where $\Phi_{t}:=\pi_{2}\circ\Phi|_{{\gamma^{*}\oplus_{1}^{N}\xi}(t)}$ and $\Psi_{t}:=\pi_{2}\circ\Psi|_{{\gamma^{*}\oplus_{1}^{N}\xi}(t)}$ with $\pi_{2}:S^{1}\times\mathbb{C}^{nN}\to\mathbb{C}^{nN}$ being the projection onto the second factor; cf. [5, 40]. Notice that this trivialization is closed under iterations, that is, the trivialization induced on $\gamma^{j}$ coincides, up to homotopy, with the $j$-th iterate of the trivialization over $\gamma$.
@@
The previous discussion allows us to define the mean index for all finite segments of Reeb orbits, not necessarily closed; see, e.g., [22]. This index depends continuously on the initial condition and the contact form (in the $C^{2}$-topology), and for closed Reeb orbits it agrees with the standard mean index. In this way, it is easy to see that the proof of [27, Lemma 3.3] works verbatim under the assumption that $c_{1}(\xi)$ is torsion. ∎
@@ 2.3. Resonance relations
The limit above exists and is rational; see [23]. When $\gamma$ is strongly non-degenerate, we have
$\hat{\chi}(\gamma)=\begin{cases}(-1)^{{\mu}(\gamma)}\ \text{ if }\gamma^{2}\text{ is good}\\ (-1)^{{\mu}(\gamma)}/2\ \text{ if }\gamma^{2}\text{ is bad}.\end{cases}$
@@
$\sum_{i=1}^{r}\frac{\hat{\chi}(\gamma_{i})}{{\hat{\mu}}(\gamma_{i})}=\chi_{\pm}(M),$ (2.3)
@@ 3. Index recurrence
A crucial ingredient in the proof of Theorem 1.6 is the following combinatorial result addressing the index behavior under iterations taken from [27, Theorem 4.1]. This result can also be deduced from the so-called enhanced common index jump theorem due to Duan, Long and Wang [19]; see also [35, 36].
@@
4. Proof of Theorem 1.6
@@ 4.1. Outline of the proof
[…]
@@
#$\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k^{\prime}_{i}})>d^{\prime}\}=\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d\},$
@@
#$\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d\}+\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})>d\}=r_{B}.$
[…]
#$\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d\text{ or }{\mu}(\gamma_{i}^{k_{i}})>d\}=r.$
@@ 4.2. Proof of the theorem
${\mathrm{HC}}^{0}_{k}(M)\cong{\mathrm{CC}}^{0}_{k}(\alpha)\cong\oplus_{\gamma\in{\mathscr{P}}^{0}(\alpha);\,{\mu}(\gamma)=k}\Lambda$ (4.1)
@@
In what follows, we will prove Theorem 1.6. For a clear exposition, we will break down the proof in claims and steps.
@@
As the reader can easily check, the only point in the proof below that uses that $M$ is a prequantization is the computation of ${\mathrm{HC}}^{0}_{*}(M)$ given by (2.1). The proof works verbatim for any closed contact manifold $M$ (under the assumption (F) or that $\alpha$ is index-admissible) whenever we have (2.1) with $c_{B}>n/2$. It will be crucial in the proof of Theorem 1.13 established in Section 5.
@@
This part of the proof follows from the work of Gürel [32, Theorem 1.5]. For the sake of completeness, we will reproduce her argument. Define
@@
The key point is the following combinatorial lemma proved in [32]:
@@
[32, Lemma 3.2] Assume that $\alpha$ has a collection of $m$ geometrically distinct contractible periodic orbits $\gamma_{1},\dotsc,\gamma_{m}$. Then for every sufficiently small $\epsilon>0$, there exist infinitely many distinct intervals $I$ of length $2n+\epsilon$ such that for some positive integers $k_{1}\geq 1,\dots,k_{m}\geq 1$ (depending on the interval), the iterated orbits $\gamma_{1}^{k_{1}},\dotsc,\gamma_{m}^{k_{m}}$ all have indexes in the interval $I$.
Since ${k_{\text{m}in}}\geq 1$, it follows from (4.1) and the previous lemma that $\alpha$ has at most $b$ simple contractible closed orbits.
@@
Note that the dimension of $M$ in [32] is $12n-$ 1 and here is $12n+$ 1.
@@
${\hat{\mu}}(\gamma_{i})>0\quad\forall i$ (4.3)
and
$\quad{\mu}(\gamma_{i}^{j})\geq{k_{\text{m}in}}\geq 1\quad\forall i,j.$ (4.4)
The inequality (4.3) holds because, since $|j{\hat{\mu}}(\gamma_{i})-{\mu}(\gamma_{i}^{j})|<n$ for every $j$, if we have some $\gamma_{i}$ such that ${\hat{\mu}}(\gamma_{i})\leq 0$ then we have either that ${\mathrm{HC}}^{0}_{k}(M)$ is non-trivial in arbitrarily large negative degrees $k$ (if ${\hat{\mu}}(\gamma_{i})<0$) or ${\mathrm{HC}}^{0}_{k}(M)$ has infinite dimension in some degree $k\in(-n,n)$ (if ${\hat{\mu}}(\gamma_{i})=0$) and both situations are impossible by (2.1).
@@
$(d,k_{1},\dots,k_{r}):=(d^{+}_{1},k^{+}_{11},\dots,k^{+}_{r1})\text{ and }(d^{\prime},k^{\prime}_{1},\dots,k^{\prime}_{r}):=(d^{-}_{1},k^{-}_{11},\dots,k^{-}_{r1}).$ (4.5)
The following lemma is one of the key steps in the proof and it is proved using the resonance relations mentioned in Section 2.3; cf. [3, Sublemma 5.2] and [27, Lemma 5.1].
@@
${\mu}(\gamma_{i}^{k_{i}-\ell})\leq d-1\quad\forall 1\leq\ell<k_{i}$ (4.6)
@@
${\mu}(\gamma_{i}^{k_{i}+\ell})\geq d+1\quad\forall\ell\geq 1.$ (4.7)
@@
#$\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k^{\prime}_{i}})>d^{\prime}\}=\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d\}.$
@@
#$\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d\}+\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})>d\}=r_{B}.$
[…]
#$\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d\text{ or }{\mu}(\gamma_{i}^{k_{i}})>d\}=r.$
@@
${\mu}(\gamma_{i}^{k_{i}-\ell})\leq d-2\quad\forall 1\leq\ell<k_{i}$ (4.8)
@@
${\mu}(\gamma_{i}^{k_{i}+\ell})\geq d+2\quad\forall\ell\geq 1.$ (4.9)
@@
#$\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k^{\prime}_{i}})>d^{\prime}+1\}=\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d-1\}.$
@@
#$\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})=d\}=\dim{\mathrm{H}}_{n}(B;\mathbb{Q}).$

### v3

@@ 1.2. Organization of the paper
The rest of this paper is organized as follows. Sections 2 and 3 furnish the necessary background for the proof of Theorem 1.6. More precisely, in Section 2.1 we briefly review several facts about positive equivariant symplectic homology. This homology is computed for suitable prequantization $S^{1}$-bundles in Section 2.2. The resonance relations for equivariant symplectic homology are presented in Section 2.3. Section 3 states the index recurrence theorem, which is a key tool in the proof of Theorem 1.6. The proof of Theorem 1.6 is established in Section 4 under the assumption that $c_{B}>n/2$. This hypothesis is dropped in Appendix A. The proof of Theorem 1.16 is presented in Section 5. Section 6 establishes results about the multiplicity of periodic Reeb orbits on lens spaces and their unit cosphere bundles, extending some results from [30]. Finally, Section 7 poses some questions concerning contact forms with finitely many closed orbits.
@@ 2.1. Equivariant symplectic homology
and an action filtration introduced by McLean and Ritter [44]; cf. [1, Section 2]. Let $\alpha$ be a non-degenerate contact form on $M$ supporting the contact structure $\xi$. Recall that a periodic orbit $\gamma$ of $\alpha$ is good if its index has the same parity of the index of the underlying simple closed orbit. Then the positive equivariant symplectic homology ${\mathrm{SH}}_{*}^{S^{1},+}(W)$ with coefficients in $\Lambda$} is the homology of a complex ${\mathrm{CC}}_{*}(\alpha)$ generated by the good closed Reeb orbits of $\alpha$; see [27, Proposition 3.3]. (More precisely, [27, Proposition 3.3] is proved assuming that $\Omega|_{\pi_{2}(W)}=0$ and using $\mathbb{Q}$-coefficients but its proof can be readily adapted to our context since it is purely algebraic; cf. [1, Section 2].) This complex is graded by the Conley–Zehnder index and filtered by the action. Furthermore, once we fix a free homotopy class of loops in $W$, the part of ${\mathrm{CC}}_{*}(\alpha)$ generated by closed Reeb orbits in that class is a subcomplex. As a consequence, the entire complex ${\mathrm{CC}}_{*}(\alpha)$ breaks down into a direct sum of such subcomplexes indexed by free homotopy classes of loops in $W$.
@@
This description of the positive equivariant symplectic homology as the homology of ${\mathrm{CC}}_{*}(\alpha)$ is not quite standard, but it is most suitable for our purposes. (We refer the reader to [27] for more details and further references and to [9, 10, 11, 12, 13, 46] for the original construction of the equivariant symplectic homology.) To see why ${\mathrm{HC}}_{*}(M):={\mathrm{SH}}_{*}^{S^{1},+}(W)$ can be obtained as the homology of a single complex generated by good closed Reeb orbits, let us first consider an admissible Hamiltonian $H$ on the symplectic completion of $W$ and focus on the orbits of $H$ with positive action. Such orbits are in a one-to-one correspondence with closed Reeb orbits $\gamma$ with action below a certain threshold $T$ depending on the slope of $H$. The $S^{1}$-equivariant Floer homology of $H$ is the homology of a Floer-type complex obtained from a non-degenerate parametrized perturbation of $H$; [12, 46]. This complex is filtered by the action. (Here we are using the action filtration introduced by McLean-Ritter [44].) The $E^{1}$-term of the resulting spectral sequence (over $\Lambda$}) is generated by the good Reeb orbits of $\alpha$ with action below $T$. Now we can (canonically, once the generators are fixed) reassemble the differentials $\partial_{r}$ into a single differential $\partial$ on ${\mathrm{CC}}_{*}(H):=E^{1}_{*,*}$ in such a way the the homology of the resulting complex is $E^{\infty}={\mathrm{HF}}_{*}^{S^{1},+}(H)$. Roughly speaking, $\partial=\partial_{1}+\partial_{2}+\ldots$, where $\partial_{r}$ is suitably “extended” from $E^{r}$ to $E^{1}$. Moreover, this procedure respects the action filtration and is functorial with respect to continuation maps. Passing to the limit in $H$, we obtain the complex ${\mathrm{CC}}_{*}(\alpha)$ as the limit of the complexes ${\mathrm{CC}}_{*}(H)$; see [27, Sections 2.5 and 3] for further details.
A remarkable observation by Bourgeois and Oancea in [13, Section 4.1.2] is that under suitable additional assumptions on the indices of closed Reeb orbits the positive equivariant symplectic homology is defined even when $M$ does not have a symplectic filling. To be more precise, we assume that $c_{1}(\xi)|_{\pi_{2}(M)}=0$ and let $\alpha$ be a non-degenerate contact form on $M$ such that all of its closed contractible Reeb orbits have Conley–Zehnder index strictly greater than 3n3-n. Furthermore, under this assumption the proof of [27, Proposition 3.3] carries over essentially word-for-word, and hence again the positive equivariant symplectic homology of $M$ can be described as the homology of a complex ${\mathrm{CC}}_{*}(\alpha)$ generated by good closed Reeb orbits of $\alpha$, graded by the Conley–Zehnder index and filtered by the action. The complex breaks down into the direct sum of subcomplexes indexed by free homotopy classes of loops in $M$. As in the fillable case, we will use the notation ${\mathrm{HC}}_{*}(M)$ and ${\mathrm{HC}}_{*}^{0}(M)$. Note that in general, in spite of the notation, this homology has slightly different properties (and hypothetically could be different) from the homology defined via a filling. For instance, it has a decomposition by the free homotopy classes of loops in $M$, but not in $W$ as when $M$ is fillable. (Intuitively, one can think of the resulting homology as defined by using a non-compact filling of $M$ by the bounded part $M\times(0,\,1]$ of the symplectization of $(M,\xi)$.)
@@ 2.2. Equivariant symplectic homology of prequantizations
The next proposition, essentially taken from [30, Proposition 3.1], shows how to compute the equivariant symplectic homology of a suitable prequantization in terms of the homology of the basis. This computation will be crucial throughout this work.
@@
    (2.1) ${\mathrm{HC}}^{0}_{\ast}(M)\cong\bigoplus_{m\in\mathbb{N}}{\mathrm{H}}_{\ast-2mc_{B}+n}(B;{\Lambda}).$
@@
    (2.2) ${\mathrm{HC}}^{0}_{\ast}(M)\cong\bigoplus_{m\in\mathbb{N}}{\mathrm{H}}_{\ast+2mc_{B}-n}(B;{\Lambda}).$
@@
    Alternatively, assume that $B$ is spherically positive monotone with $c_{B}\geq 2$ and, as in the hypothesis (NF) of Theorem 1.6, $c_{1}(\xi)|_{H_{2}(M,\mathbb{Q})}=0$ and $\alpha$ is a non-degenerate contact form on $(M,\xi)$ such that all contractible closed Reeb orbits have index greater than 3n3-n. Then (2.1) also holds.
@@
First note that, by the universal coefficient theorem, ${\mathrm{H}}_{*}(B;{\Lambda})\cong{\mathrm{H}}_{*}(B;\mathbb{Q})\otimes{\Lambda}$ since $\Lambda$} is a field. Then the proof goes word-for-word the proof of [30, Proposition 3.1] using the Novikov field and the action filtration mentioned in the previous section, except in the proof of [30, Lemma 3.3], used in the proof of item (b), where it is assumed that $c_{1}(\xi)=0$ while here we are allowing $c_{1}(\xi)$ to be torsion. This lemma establishes that a sufficiently small non-degenerate perturbation of the connection form is index-admissible, that is, every contractible periodic orbit $\gamma$ satisfies ${\mu}(\gamma)>3-n$. The assumption in [30, Lemma 3.3] that $c_{1}(\xi)=0$ is used to ensure that the determinant line bundle $\Lambda^{n}_{\mathbb{C}}\xi$ is trivial. When $c_{1}(\xi)$ is torsion it is no longer true in general, but we have that $(\Lambda_{\mathbb{C}}^{n}\xi)^{\otimes N}$ is a trivial line bundle, where $N$ is the smallest positive integer such that $Nc_{1}(\xi)=0$. Choose a trivialization $\tau:(\Lambda_{\mathbb{C}}^{n}\xi)^{\otimes N}\to M\times\mathbb{C}$ which corresponds to a choice of a non-vanishing section $\mathfrak{s}$} of $(\Lambda_{\mathbb{C}}^{n}\xi)^{\otimes N}$. The choice of this trivialization furnishes a unique way to symplectically trivialize $\oplus_{1}^{N}\xi$ along periodic orbits of $\alpha$ up to homotopy. As a matter of fact, given a periodic orbit $\gamma$, let $\Phi:{\gamma^{*}\oplus_{1}^{N}\xi}\to S^{1}\times\mathbb{C}^{nN}$ be a trivialization of $\oplus_{1}^{N}\xi$ over $\gamma$ as a Hermitian vector bundle such that its highest complex exterior power coincides with $\tau$. This condition fixes the homotopy class of $\Phi$: given any other such trivialization $\Psi$ we have, for every $t\in S^{1}$, that $\Phi_{t}\circ\Psi_{t}^{-1}:\mathbb{C}^{nN}\to\mathbb{C}^{nN}$ has complex determinant equal to one and therefore the Maslov index of the symplectic path $t\mapsto\Phi_{t}\circ\Psi_{t}^{-1}$ vanishes, where $\Phi_{t}:=\pi_{2}\circ\Phi|_{{\gamma^{*}\oplus_{1}^{N}\xi}(t)}$ and $\Psi_{t}:=\pi_{2}\circ\Psi|_{{\gamma^{*}\oplus_{1}^{N}\xi}(t)}$ with $\pi_{2}:S^{1}\times\mathbb{C}^{nN}\to\mathbb{C}^{nN}$ being the projection onto the second factor; cf. [5, 43]. Notice that this trivialization is closed under iterations, that is, the trivialization induced on $\gamma^{j}$ coincides, up to homotopy, with the $j$-th iterate of the trivialization over $\gamma$.
@@
The previous discussion allows us to define the mean index for all finite segments of Reeb orbits, not necessarily closed; see, e.g., [25]. This index depends continuously on the initial condition and the contact form (in the $C^{2}$-topology), and for closed Reeb orbits it agrees with the standard mean index. In this way, it is easy to see that the proof of [30, Lemma 3.3] works verbatim under the assumption that $c_{1}(\xi)$ is torsion. ∎
@@ 2.3. Resonance relations
The limit above exists and is rational; see [26]. When $\gamma$ is strongly non-degenerate, i.e. all iterates of $\gamma$ are non-degenerate, we have
[…]
@@
(2.3) $\sum_{i=1}^{r}\frac{\hat{\chi}(\gamma_{i})}{{\hat{\mu}}(\gamma_{i})}=\chi_{\pm}(M),$
@@ 3. Index recurrence
A crucial ingredient in the proof of Theorem 1.6 is the following combinatorial result addressing the index behavior under iterations taken from [30, Theorem 4.1]. This result can also be deduced from the so-called enhanced common index jump theorem due to Duan, Long and Wang [22]; see also [38, 39].
@@
4. Proof of Theorem 1.6 under the assumption that $c_{B}>n/2$
Throughout this section, we will assume that $c_{B}>n/2$. As explained in the introduction, this hypothesis is dropped in Appendix A.
@@ 4.1. Outline of the proof
So we have finitely many simple contractible orbits $\{\gamma_{1},\dots,\gamma_{r}\}$ and we have to show that $r=r_{B}$. Since the differential in ${\mathrm{CC}}_{*}^{0}(\alpha)$ vanishes, we have, using (2.1), that every contractible orbit $\gamma$ satisfies ${\mu}(\gamma)=n\ \text{(mod 2)}$, ${\mu}(\gamma)\geq{k_{\text{m}in}}:=\min\{k\in\mathbb{Z};\ {\mathrm{HC}}^{0}_{k}(M)\neq 0\}\geq 1$ and ${\hat{\mu}}(\gamma)>0$. Using this and Theorem 3.1, we find integers $d,k_{1},\dots,k_{r}$ such that $d=2sc_{B}$ for some $s\in\mathbb{N}$,
@@
$\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k^{\prime}_{i}})>d^{\prime}\}=\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d\},$
@@
$\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d\}+\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})>d\}=r_{B}.$
But $d=2sc_{B}$ is even which implies that ${\mu}(\gamma_{i}^{k_{i}})\neq d$ since ${\mu}(\gamma_{i}^{k_{i}})=n\ \text{(mod 2)}$ is odd. Therefore,
$\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d\text{ or }{\mu}(\gamma_{i}^{k_{i}})>d\}=r.$
@@ 4.2. Proof of the theorem
(4.1) ${\mathrm{HC}}^{0}_{k}(M)\cong{\mathrm{CC}}^{0}_{k}(\alpha)\cong\oplus_{\gamma\in{\mathscr{P}}^{0}(\alpha);\,{\mu}(\gamma)=k}\Lambda$
@@
It follows from (4.1) and (2.1) that, for every contractible orbit $\gamma$ of $\alpha$,
(4.2) ${\hat{\mu}}(\gamma)>0.$
Indeed, since $|j{\hat{\mu}}(\gamma)-{\mu}(\gamma^{j})|<n$ for every $j$, if we have some $\gamma$ such that ${\hat{\mu}}(\gamma)\leq 0$ then we have either that ${\mathrm{HC}}^{0}_{k}(M)$ is non-trivial in arbitrarily large negative degrees $k$ (if ${\hat{\mu}}(\gamma)<0$) or ${\mathrm{HC}}^{0}_{k}(M)$ has infinite dimension in some degree $k\in(-n,n)$ (if ${\hat{\mu}}(\gamma)=0$) and both situations are impossible by (2.1).
@@
In what follows, we will prove Theorem 1.6 under the assumption that $c_{B}>n/2$. For a clear exposition, we will break down the proof in claims and steps.
@@
As the reader can easily check, the only point in the proof below that uses that $M$ is a prequantization is the computation of ${\mathrm{HC}}^{0}_{*}(M)$ given by (2.1). The proof works verbatim for any closed contact manifold $M$ (under the assumption (F) or that $\alpha$ is index-admissible) whenever we have (2.1) with $c_{B}>n/2$. (This assumption on $c_{B}$ can be completely dropped as explained in Appendix A.) It will be crucial in the proof of Theorem 1.16 established in Section 5.
@@
This part of the proof follows from the work of Gürel [35, Theorem 1.5]. For the sake of completeness, we will reproduce her argument. Define
@@
The key point is the following combinatorial lemma proved in [35] which uses the aforementioned fact that every contractible orbit of $\alpha$ has positive mean index:
@@
[35, Lemma 3.2] Assume that $\alpha$ has a collection of $m$ geometrically distinct contractible periodic orbits $\gamma_{1},\dotsc,\gamma_{m}$. Then for every sufficiently small $\epsilon>0$, there exist infinitely many disjoint intervals $I$ of length $2n+\epsilon$ such that for some positive integers $k_{1}\geq 1,\dots,k_{m}\geq 1$ (depending on the interval), the iterated orbits $\gamma_{1}^{k_{1}},\dotsc,\gamma_{m}^{k_{m}}$ all have indexes in the interval $I$.
Since $k_{\text{m}in}$$\infty$, it follows from (4.1) and the previous lemma that $\alpha$ has at most $b$ simple contractible closed orbits.
@@
Note that the dimension of $M$ in [35] is 2n12n-1 and here is $2n+1$.
@@
(4.4) $\quad{\mu}(\gamma_{i}^{j})\geq{k_{\text{m}in}}\geq 1\quad\forall i,j.$
@@
[…]
The following lemma is one of the key steps in the proof and it is proved using the resonance relations mentioned in Section 2.3; cf. [3, Sublemma 5.2] and [30, Lemma 5.1].
@@
(4.6) ${\mu}(\gamma_{i}^{k_{i}-\ell})\leq d-1\quad\forall 1\leq\ell<k_{i}$
@@
(4.7) ${\mu}(\gamma_{i}^{k_{i}+\ell})\geq d+1\quad\forall\ell\geq 1.$
@@
$\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k^{\prime}_{i}})>d^{\prime}\}=\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d\}.$
@@
$\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d\}+\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})>d\}=r_{B}.$
But $d=2sc_{B}$ is even which implies that ${\mu}(\gamma_{i}^{k_{i}})\neq d$ since ${\mu}(\gamma_{i}^{k_{i}})=n\ \text{(mod 2)}$ is odd. Therefore,
$\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d\text{ or }{\mu}(\gamma_{i}^{k_{i}})>d\}=r.$
@@
(4.10) ${\mu}(\gamma_{i}^{k_{i}-\ell})\leq d-2\quad\forall 1\leq\ell<k_{i}$
@@
(4.11) ${\mu}(\gamma_{i}^{k_{i}+\ell})\geq d+2\quad\forall\ell\geq 1.$
@@
$\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k^{\prime}_{i}})>d^{\prime}+1\}=\#\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})<d-1\}.$
@@
$\{1\leq i\leq r;\ {\mu}(\gamma_{i}^{k_{i}})=d\}=\dim{\mathrm{H}}_{n}(B;\mathbb{Q}).$